OpenStudy (anonymous):

(a^3b^12c^2)*(a^5c^2)/(b^5c^4)^0= A. a^14b^15c^9 B. a^8b^12c^4 C. a^8b^17c^8 D. a^15b^12c^4

OpenStudy (mathlegend):

Well in the denominator we know that anything to the 0 power is 1 so you can just eliminate that completely.

OpenStudy (mathlegend):

Now, we have left.... (a^3b^12c^2)*(a^5c^2) When you multiply we add exponents

OpenStudy (mathlegend):

a^3*a^5 = a^8 b^12 c^2*c^2= c^4 Answer is... B

OpenStudy (mathlegend):

@haleylou56 do you understand how I did that?

OpenStudy (anonymous):

yes i did, thank you

OpenStudy (mathlegend):

:)